# High-Dimensional Bayesian Model Comparison in Cosmology with GPU-accelerated Nested Sampling and Neural Emulators

### v1

@@ 1 Introduction
Bayesian inference offers a principled statistical framework for parameter estimation and model comparison, and is widely employed in astrophysics for distinguishing between competing cosmological theories (Trotta, 2008). The advent of large-scale surveys such as Euclid (Euclid Collaboration et al., 2025) [^1] https://www.euclid-ec.org/, the Vera C. Rubin Observatory (Ivezić et al., 2019) [^2] https://www.lsst.org/, and the Nancy Grace Roman Space Telescope (Spergel et al., 2015) [^3] https://roman.gsfc.nasa.gov/, alongside increasingly sophisticated theoretical models, has led to a significant increase in both data volume and model dimensionality. While parameter estimation remains manageable in this new scale of problem, the accurate computation of the Bayesian evidence stands as a distinct computational hurdle. Bayesian evidence is very valuable as a model comparison tool, as it weighs goodness of fit against model complexity, naturally penalizing over-fit models (Trotta, 2007; Lovick et al., 2024).
@@
To address these high-dimensional inference challenges, the field has undergone significant algorithmic innovation. Advanced Markov chain Monte Carlo (MCMC) techniques, including gradient-based samplers like Hamiltonian Monte Carlo (HMC; Duane et al., 1987; Neal, 1996) and the No-U-Turn Sampler (NUTS; Hoffman and Gelman, 2014) which utilise differentiability for efficiency, Variational Inference (VI) offering an optimization-based alternative (Blei et al., 2017; Hoffman et al., 2013; Uzsoy et al., 2024), and machine learning-augmented approaches, such as emulators (e.g. Albers et al., 2019; Manrique-Yus and Sellentin, 2019; Mootoovaloo et al., 2020; Arico’ et al., 2022; Mootoovaloo et al., 2022; Spurio Mancini et al., 2022; Piras and Spurio Mancini, 2023; El Gammal et al., 2023; Günther, 2024; Bonici et al., 2024, 2024; Günther et al., 2025) and the learned harmonic mean estimator for decoupled Bayesian evidence estimation (McEwen et al., 2021; Polanska et al., 2024, 2025; Lin et al., 2025), have emerged as promising methods. Concurrently, the parallel processing capabilities of Graphics Processing Units (GPUs) have been identified as an important development, accelerating likelihood evaluations and sample generation across diverse methodologies (Gu et al., 2022; García-Quirós et al., 2025; Prathaban et al., 2025), and extending the scope of feasible analyses.
@@
Nested Sampling (NS; Skilling, 2006) is a notable meta-algorithm that offers simultaneous parameter inference and Bayesian evidence calculation, a key advantage for rigorous model selection, while also effectively exploring complex, multimodal posterior landscapes. Although traditional CPU-based NS implementations have encountered scalability limitations in high-dimensional settings (Feroz et al., 2009), recent work to adapt NS to efficiently utilize GPU-hardware (Yallup et al., 2025) has pushed the capabilities of the algorithm much further than its CPU counterpart. Our work demonstrates that GPU-accelerated NS is significantly faster than its CPU counterparts on various cosmological likelihoods, while maintaining accurate posterior inference. This paper is structured as follows. In Section 2 we describe in high-level the sampling algorithm and likelihoods demonstrated in this work, in section 3 we display the model evidences and computation times of analyses, then in section 4 we discuss and unpack the relative performance of the analyses, as well as the advantages and disadvantages between nested sampling and HMC with decouple evidence calculation.
@@ 2.1 Nested Sampling
Nested Sampling (NS; Skilling, 2006) is a Monte Carlo method designed to draw posterior samples and calculate the Bayesian evidence, $\mathcal{Z}$, defined as the integral of the likelihood $\mathcal{L}$ over the prior distribution $\pi$ of the parameters $\theta$:
@@
The algorithm iteratively replaces the point with the lowest likelihood among a set of “live points” with a new point drawn from the prior, constrained to have a likelihood higher than the point being replaced. This process allows the evidence to be accumulated as a sum of likelihood values weighted by successively smaller prior volumes. NS is well established as one of the primary methods for estimating the Bayesian evidence of models in cosmology, appearing in many contemporary high profile results (e.g. Planck Collaboration et al., 2020; Abbott et al., 2022; Wright et al., 2025). For a full review of NS aimed at physical scientists, see Ashton et al. 2022.
The key challenge of implementing NS is to efficiently draw samples from the prior within these progressively shrinking, hard likelihood boundaries. Slice sampling (Neal, 2003) is an MCMC technique well-suited for this task, as it can adaptively sample from such constrained distributions without requiring manual tuning of proposal scales. Usage of slice sampling within nested sampling was popularised in Handley et al. (2015, 2015). Slice sampling works by sampling uniformly from an auxiliary variable defining a “slice” under the likelihood surface and then sampling the parameters from the prior surface subject to a hard likelihood constraint. The work of Yallup et al. (2025) provided a generic implementation of the nested sampling algorithm, as well as a tailored implementation of a slice sampling based algorithm that was designed specifically to be amenable to the massive parallelism opportunities of modern GPUs. The latter implementation is styled as Nested Slice Sampling (NSS), and we adopt this specific algorithm in this work.
@@ 2.2 Parallel sampling and the future of inference
There are two main mechanisms by which a sampling code can be accelerated on GPU hardware. Firstly by leveraging the massively parallel SIMD (Same Instruction Multiple Data) paradigm to distribute a calculation over many threads, and secondly by utilising gradient information to improve the mixing of the algorithm itself. The frameworks used to encode calculations on GPUs typically afford fast evaluation of gradients via automatic differentiation (Bradbury et al., 2018), so this information is typically readily available for use. GPU native inference paradigms, such as momentum gradient descent as used to train neural networks (Kingma and Ba, 2017), are so successful as they utilize both of these mechanisms. Leveraging both these mechanisms in scientific probabilistic inference codes optimally is more challenging, and is the focus of this work.
@@
Hamiltonian Monte Carlo is a common choice in many scientific analyses, as a high level algorithm this can exploit the fast evaluation of gradients for rapid convergence. Popular self-tuning variants of HMC, such as NUTS, have been established to work well on many Bayesian inference problems across a variety of scales. However the large variation in walk length incurred in these schemes makes parallelisation of multiple chains much less profitable, with contemporary work seeking to derive robust alternatives that are more amenable to GPU hardware (Hoffman and Sountsov, 2022; Hoffman et al., 2021).
@@
Nested sampling, despite its prominence in many cosmological applications, does not make widespread use of either gradients or SIMD parallelism in its popular implementations (Handley et al., 2015; Buchner, 2021; Speagle, 2020; Albert, 2020). Realising the full potential of gradients within the nested sampling algorithm for general problems is an active area of research (Betancourt et al., 2011; Lemos et al., 2023; Cai et al., 2022) we leave for future investigation, noting that establishing benchmark sampling problems in differentiable frameworks is a useful by-product of this work. Nested sampling can be parallelised by selecting $k$ live points with the lowest likelihoods rather than just the lowest, and evolving them with the likelihood constraint in parallel. This adaptation does not introduce new approximations into the core NS algorithm itself; rather, it significantly enhances computational speed by harnessing parallel hardware capabilities. If the generation of new points can be vectorized or batched, this represents a substantial potential speed-up for GPU execution, and evolving many points at once should scale as $\mathcal{O}(1)$ with the number of points, up to the parallelisation limit of the GPU. In practice Yallup et al. (2025) demonstrated that perfect scaling is hard to achieve, however we demonstrate in this work that it is possible to gain orders of magnitude speed-up over classic NS pipelines.
@@
The primary computational bottleneck in this parallel scheme is the speed of bulk likelihood evaluations. This motivates the use of jax-based likelihoods and end-to-end cosmological pipelines (e.g. Campagne et al., 2023; Ruiz-Zapatero et al., 2024; Balkenhol et al., 2024). Further acceleration can be obtained by replacing traditional Boltzmann solvers such as CAMB (Lewis and Challinor, 2011) or CLASS (Blas et al., 2011) with differentiable implementations (e.g. Hahn et al., 2024), which are designed to run efficiently on modern hardware. Ultimately, the largest speed-ups are achieved through emulation with neural-network surrogates such as CosmoPower-JAX (Piras and Spurio Mancini, 2023), a JAX-based extension of the CosmoPower emulator framework (Spurio Mancini et al., 2022), which fully exploits the parallel processing capabilities of GPUs. The NSS implementation is provided in the blackjax (Cabezas et al., 2024) framework. As a native jax (Babuschkin et al., 2020) compilable sampling framework, this is well positioned to exploit the parallelisation opportunities present at the likelihood level, reducing key factors such as communication overhead between devices that can plague hardware accelerated workflows.
@@
In contrast, MCMC-based methods can access GPU vectorization by running multiple independent chains in parallel. The generation of a new sample within a single chain is an inherently sequential process, so its speed is limited by the evaluation time of the likelihood function. While running many chains in parallel is an effective use of the GPU, this strategy requires that each chain independently reaches convergence, so the need for a “burn-in” period for each chain to discard its initial non-stationary samples (Hoffman and Gelman, 2014) means that the efficiency does not scale as directly as batching an NS run. As the target of our analysis is performing model comparison, on top of using HMC (NUTS) to draw samples, we employ the harmonic framework (McEwen et al., 2021), which trains a normalising flow on posterior samples in order to compute the evidence via the learned harmonic mean estimator. We utilise the same flow architecture and training procedure as in (Piras et al., 2024), training the flow on 30% of chains and using the remaining 70% for inference to compute the evidence. In this work polychord (Handley et al., 2015, 2015) provides the comparison for a CPU-based nested sampler.
@@
Whilst the NS analyses we develop in this work will prove to be scalable and competitive with other start-of-the-art model comparison techniques, without use of gradients it is important to recognise that full field level inference in cosmology (e.g. Jasche and Wandelt, 2013; Porqueres et al., 2021a, b; Bayer et al., 2023; Spurio Mancini et al., 2024; Babić et al., 2025; Lanzieri et al., 2025) is beyond the scope of this current NS implementation. It is however important to explore limitations of techniques, and developing competitive strategies in consistent frameworks (free from potential constraints of legacy codes), is vital to pushing the frontiers of inference problems in the physical sciences.
@@
Table 1: Prior distributions and fiducial values for the parameters of the cosmic shear model, split into cosmological, Baryonic, and nuisance parameters. All priors used are uniform ($\mathcal{U}$) or Gaussian ($\mathcal{N}$). All values and ranges in this table follow Piras and Spurio Mancini (2023).
@@
$\omega_{b}=\Omega_{b}h^{2}$ $\mathcal{U}(0.01875,0.02625)$ 0.02242
@@ 2.3 CMB and Cosmic Shear Analyses
We demonstrate the sampler on two cosmological problems: a cosmic variance-limited CMB power spectrum analysis and a cosmic shear analysis. The CMB analysis, a standard 6-dimensional problem, functions as a benchmark and demonstration of the sampler’s speed when used on a highly vectorisable likelihood, and its accuracy on a foundational application (Planck Collaboration et al., 2020). The Cosmic Shear analysis, with 7 cosmological and 30 nuisance parameters, presents a higher dimensional challenge that represents the edge of what CPU-based nested sampling can manage, and offers a direct comparison to the HMC + learned Harmonic Mean Estimator analysis of Piras et al. (2024). Figure 1 shows the vectorisation behaviour of both of these likelihoods, namely how they scale in execution time when called with batches of parameters at a time.
(a) CMB Likelihood as timed on an L4 GPU, with near-perfect vectorization up to $\mathcal{O}(10^{3})$ parallel calls.
(b) Shear Likelihood timed on an A100 GPU, where using interpolation introduces some vectorisation.
Figure 1: Execution time for batched likelihood evaluations for both the CMB and Cosmic Shear likelihoods. Here the y-axis refers to the time relative to calling the likelihood on a single set of parameters.

### v2

@@ 1 Introduction
Bayesian inference offers a principled statistical framework for parameter estimation and model comparison, and is widely employed in astrophysics for distinguishing between competing cosmological theories (Trotta,, 2008). The advent of large-scale surveys such as Euclid (Euclid Collaboration et al.,, 2025) [^1] https://www.euclid-ec.org/, the Vera C. Rubin Observatory (Ivezić et al.,, 2019) [^2] https://www.lsst.org/, and the Nancy Grace Roman Space Telescope (Spergel et al.,, 2015) [^3] https://roman.gsfc.nasa.gov/, alongside increasingly sophisticated theoretical models, has led to a significant increase in both data volume and model dimensionality. While parameter estimation remains manageable in this new scale of problem, the accurate computation of the Bayesian evidence stands as a distinct computational hurdle. Bayesian evidence is very valuable as a model comparison tool, as it weighs goodness of fit against model complexity, naturally penalising over-fit models (Trotta,, 2007; Lovick et al.,, 2024).
@@
To address these high-dimensional inference challenges, the field has undergone significant algorithmic innovation. Advanced Markov chain Monte Carlo (MCMC) techniques, including gradient-based samplers like Hamiltonian Monte Carlo (HMC; Duane et al.,, 1987; Neal,, 1996) and the No-U-Turn Sampler (NUTS; Hoffman and Gelman,, 2014) which utilise differentiability for efficiency, Variational Inference (VI) offering an optimisation-based alternative (Hoffman et al.,, 2013; Blei et al.,, 2017; Uzsoy et al.,, 2024), and machine learning-augmented approaches, such as emulators (e.g. Albers et al.,, 2019; Manrique-Yus and Sellentin,, 2019; Mootoovaloo et al.,, 2020; Arico’ et al.,, 2022; Mootoovaloo et al.,, 2022; Spurio Mancini et al.,, 2022; Piras and Spurio Mancini,, 2023; El Gammal et al.,, 2023; Günther,, 2024; Bonici et al., 2024a, ; Bonici et al., 2024b, ; Günther et al.,, 2025) and the learned harmonic mean estimator for decoupled Bayesian evidence estimation (McEwen et al.,, 2021; Polanska et al.,, 2024, 2025; Lin et al.,, 2025), have emerged as promising methods. Concurrently, the parallel processing capabilities of Graphics Processing Units (GPUs) have been identified as an important development, accelerating likelihood evaluations and sample generation across diverse methodologies (Gu et al.,, 2022; García-Quirós et al.,, 2025; Prathaban et al.,, 2025), and extending the scope of feasible analyses.
@@
Nested Sampling (NS; Skilling,, 2006) is a notable sampling framework that offers simultaneous parameter inference and Bayesian evidence calculation, a key advantage for rigorous model selection, while also effectively exploring complex, multimodal posterior landscapes. Although traditional CPU-based NS implementations have encountered scalability limitations in high-dimensional settings (Feroz et al.,, 2009), recent work to adapt NS to efficiently utilise GPU-hardware (Yallup et al.,, 2025) has pushed the capabilities of the algorithm much further than its CPU counterpart. Our work demonstrates that GPU-accelerated NS is significantly faster than its CPU counterparts on various cosmological likelihoods, while maintaining accurate posterior inference.
This paper is structured as follows. In Section 2 we describe at high-level the sampling algorithm and likelihoods demonstrated in this work, in Section 3 we display the model evidences and computation times of analyses, then in Section 4 we discuss and unpack the relative performance of the analyses, as well as the advantages and disadvantages between nested sampling and HMC with decoupled evidence calculation.
@@ 2.1 Nested Sampling
Nested Sampling (NS) is a Monte Carlo method designed to draw posterior samples and calculate the Bayesian evidence, $\mathcal{Z}$, defined as the integral of the likelihood $\mathcal{L}$ over the prior distribution $\pi$ of the parameters $\theta$:
@@
The evidence can then be used in Bayesian model comparison, where the posterior probabilities of models can be calculated using:
$\frac{P(M_{0}|\text{d})}{P(M_{1}|\text{d})}=B_{01}\frac{\pi(M_{0})}{\pi(M_{1})}\ ,\quad B_{01}=\frac{\mathcal{Z}_{0}}{\mathcal{Z}_{1}}\ ,$ (2)
where the Bayes Factor $B_{01}$ and describes the relative strength of different models as they fit the observed data (Jeffreys,, 1961).
The algorithm iteratively replaces the point with the lowest likelihood among a set of “live points” with a new point drawn from the prior, constrained to have a likelihood higher than the point being replaced. This process allows the evidence to be accumulated as a sum of likelihood values weighted by successively smaller prior volumes. NS is well established as one of the primary methods for estimating the Bayesian evidence of models in cosmology, appearing in many contemporary high profile results (e.g. Planck Collaboration et al.,, 2020; Abbott et al.,, 2022; Wright et al.,, 2025). For a full review of NS aimed at physical scientists, see Ashton et al., 2022.
The key challenge of implementing NS is to efficiently draw samples from the prior within these progressively shrinking, hard likelihood boundaries. Slice sampling (Neal,, 2003) is an MCMC technique well-suited for this task, as it can adaptively sample from such constrained distributions without requiring manual tuning of proposal scales. Usage of slice sampling within nested sampling was popularised in Handley et al., 2015a ; Handley et al., 2015b . Slice sampling works by sampling uniformly from an auxiliary variable defining a “slice” under the likelihood surface, and then sampling the parameters from the prior surface subject to a hard likelihood constraint. The work of Yallup et al., (2025) provided a generic implementation of the nested sampling algorithm, as well as a tailored implementation of a slice sampling based algorithm that was designed specifically to be amenable to the massive parallelism opportunities of modern GPUs. The latter implementation is styled as Nested Slice Sampling (NSS), and we adopt this specific algorithm in this work.
@@ 2.2 Parallel sampling and the future of inference
There are two main mechanisms by which a sampling code can be accelerated on GPU hardware. Firstly by leveraging the massively parallel SIMD (Same Instruction Multiple Data) paradigm to distribute a calculation over many threads, and secondly by utilising gradient information to improve the mixing of the algorithm itself. The frameworks used to encode calculations on GPUs typically afford fast evaluation of gradients via automatic differentiation (Bradbury et al.,, 2018), so this information is typically readily available for use. GPU native inference paradigms, such as momentum gradient descent as used to train neural networks (Kingma and Ba,, 2017), are so successful as they utilise both of these mechanisms. Leveraging both these mechanisms in scientific probabilistic inference codes optimally is more challenging, and is the focus of this work.
@@
HMC is a common choice in many scientific analyses, as a high level algorithm this can exploit the fast evaluation of gradients for rapid convergence. Popular self-tuning variants of HMC, such as NUTS, have been established to work well on many Bayesian inference problems across a variety of scales. However, the large variation in walk length incurred in these schemes makes parallelisation of multiple chains much less profitable, with contemporary work seeking to derive robust alternatives that are more amenable to GPU hardware (Hoffman et al.,, 2021; Hoffman and Sountsov,, 2022).
@@
Nested sampling, despite its prominence in many cosmological applications, does not make widespread use of either gradients or SIMD parallelism in its popular implementations (Handley et al., 2015b, ; Buchner,, 2021; Speagle,, 2020; Albert,, 2020). Realising the full potential of gradients within the nested sampling algorithm for general problems is an active area of research (Betancourt et al.,, 2011; Cai et al.,, 2022; Lemos et al.,, 2023) we leave for future investigation, noting that establishing benchmark sampling problems in differentiable frameworks is a useful by-product of this work. Nested sampling can be parallelised by selecting $k$ live points with the lowest likelihoods rather than just the lowest, and evolving them with the likelihood constraint in parallel. This adaptation does not introduce new approximations into the core NS algorithm itself; rather, it significantly enhances computational speed by harnessing parallel hardware capabilities. If the generation of new points can be vectorised or batched, this represents a substantial potential speed-up for GPU execution, and evolving many points at once should scale as $\mathcal{O}(1)$ with the number of points, up to the parallelisation limit of the GPU. In practice, Yallup et al., (2025) demonstrated that perfect scaling is hard to achieve, however we demonstrate in this work that it is possible to gain orders of magnitude speed-up over classic NS pipelines.
@@
The primary computational bottleneck in this parallel scheme is the speed of bulk likelihood evaluations. This motivates the use of JAX-based likelihoods and end-to-end cosmological pipelines (e.g. Campagne et al.,, 2023; Ruiz-Zapatero et al.,, 2024; Balkenhol et al.,, 2024; Bonici et al.,, 2025; Reymond et al.,, 2025). Further acceleration can be obtained by replacing traditional Boltzmann solvers such as CAMB (Lewis and Challinor,, 2011) or CLASS (Blas et al.,, 2011) with differentiable implementations (e.g. Hahn et al.,, 2024), which are designed to run efficiently on modern hardware. Ultimately, the largest speed-ups are achieved through emulation with neural-network surrogates such as CosmoPower-JAX (Piras and Spurio Mancini,, 2023), a JAX-based extension of the CosmoPower emulator framework (Spurio Mancini et al.,, 2022), which fully exploits the parallel processing capabilities of GPUs. The NSS implementation is provided in the blackjax (Cabezas et al.,, 2024) framework. As a native JAX (Babuschkin et al.,, 2020) compilable sampling framework, this is well positioned to exploit the parallelisation opportunities present at the likelihood level, reducing key factors such as communication overhead between devices that can plague hardware accelerated workflows.
@@
In contrast, MCMC-based methods can access GPU vectorisation by running multiple independent chains in parallel. The generation of a new sample within a single chain is an inherently sequential process, so its speed is limited by the evaluation time of the likelihood function. While running many chains in parallel is an effective use of the GPU, this strategy requires that each chain independently reaches convergence, so the need for a “burn-in” period for each chain to discard its initial non-stationary samples (Hoffman and Gelman,, 2014) means that the efficiency does not scale as directly as batching an NS run.
As the target of our analysis is performing model comparison, on top of using HMC (NUTS) to draw samples, we employ the harmonic framework (McEwen et al.,, 2021; Polanska et al.,, 2025), which trains a normalising flow on posterior samples in order to compute the evidence via the learned harmonic mean estimator. We utilise the same flow architecture and training procedure as in Piras et al., (2024), training the flow on 30% of chains and using the remaining 70% for inference to compute the evidence. In this work polychord provides the comparison for a CPU-based nested sampler.
@@
Whilst the NS analyses we develop in this work will prove to be scalable and competitive with other state-of-the-art model comparison techniques, without gradients it is important to recognise that full field-level inference in cosmology (e.g. Jasche and Wandelt,, 2013; Leclercq et al.,, 2015; Jasche and Lavaux,, 2019; Lavaux et al.,, 2019; Ramanah et al.,, 2019; Porqueres et al., 2021a, ; Porqueres et al., 2021b, ; Bayer et al.,, 2023; Andrews et al.,, 2023; Loureiro et al.,, 2023; Sellentin et al.,, 2023; Stadler et al.,, 2023; Stopyra et al.,, 2023; Spurio Mancini et al.,, 2024; Babić et al.,, 2025; Lanzieri et al.,, 2025; McAlpine et al.,, 2025; see also Leclercq,, 2025 for a recent review), is beyond the scope of this current NS implementation. It is however important to explore the limitations of current techniques, and developing competitive strategies in consistent frameworks (free from potential constraints of legacy codes) is vital to pushing the frontiers of inference problems in the physical sciences.
@@
Table 1: Prior distributions and fiducial values for the parameters of the cosmic shear model, split into cosmological, baryonic, and nuisance parameters. All priors used are uniform ($\mathcal{U}$) or Gaussian ($\mathcal{N}$). All values and ranges in this table follow Piras and Spurio Mancini, (2023).
@@
$\omega_{\rm{b}}=\omega_{\rm{b}}h^{2}$ $\mathcal{U}(0.01875,0.02625)$ 0.02242
@@ 2.3 CMB and Cosmic Shear Analyses
We demonstrate the sampler on two cosmological problems: a cosmic variance-limited CMB power spectrum analysis and a cosmic shear analysis. The CMB analysis, a standard 6-dimensional problem, functions as a benchmark and demonstration of the sampler’s speed when used on a highly vectorisable likelihood, and its accuracy on a foundational application (Planck Collaboration et al.,, 2020). The cosmic shear analysis, with 7 cosmological and 30 nuisance parameters, presents a higher dimensional challenge that represents the edge of what CPU-based nested sampling can manage, and offers a direct comparison to the HMC + learned harmonic mean estimator analysis of Piras et al., (2024). Figure 1 shows the vectorisation behaviour of both of these likelihoods, namely how they scale in execution time when called with batches of parameters at a time.
(a) CMB likelihood as timed on an L4 GPU, with near-perfect vectorisation up to approximately $10^{3}$ parallel calls.
(b) Cosmic shear likelihood timed on an A100 GPU, where using interpolation introduces some vectorisation.
Figure 1: Execution time for batched likelihood evaluations for both the CMB and cosmic shear likelihoods. Here the y-axis refers to the time relative to calling the likelihood on a single set of parameters. The interpolated likelihood involved evaluating the Matter Power Spectrum at fewer $z$-values, explained in detail in Appendix B.